# Babysit A Language Model From Scratch: Interactive Language Learning by Trials and Demonstrations

### v1

@@ Abstract
Humans are efficient language learners and inherently social creatures. Our language development is largely shaped by our social interactions, for example, the demonstration and feedback from caregivers. Contrary to human language learning, recent advancements in large language models have primarily adopted a non-interactive training paradigm, and refined pre-trained models through feedback afterward. In this work, we aim to examine how corrective feedback from interactions influences neural language acquisition from the ground up through systematically controlled experiments, assessing whether it contributes to learning efficiency in language models. We introduce a trial-and-demonstration (TnD) learning framework that incorporates three components: student trials, teacher demonstrations, and a reward conditioned on language competence at various developmental stages. Our experiments reveal that the TnD approach accelerates word acquisition for student models of equal and smaller numbers of parameters, and we highlight the significance of both trials and demonstrations. We further show that the teacher’s choices of words influence students’ word-specific learning efficiency, and a practice-makes-perfect effect is evident by a strong correlation between the frequency of words in trials and their respective learning curves. Our findings suggest that interactive language learning, with teacher demonstrations and student trials, can facilitate efficient word learning in language models.
@@ 1 Introduction
Humans are social beings and we learn language from interactions [67, 12, 41, 28, 64]. Long before children’s linguistic skills are mature, they could engage in early forms of conversational exchange with others [24, 16]. A critical component of social interactions in language acquisition is the feedback provided by the caregivers [70], for example, the communicative feedback [3, 38] that highlights the success and failure of communication. Besides, there have been extensive studies on another type of feedback that is more direct and emphasizes the responses from caregivers, which offer corrections to possible errors in children’s speech, encompassing variants like negative evidence, reformulations, or recasts, collectively referred to as corrective feedback [21, 15, 52, 26]. Unlike human learners who acquire language skills through feedback during interactions, most language models differ in terms of their inductive biases and data sources [71]. These models typically learn from massive text corpora using cross-entropy loss for self-supervised learning.
Recently, several lines of cognitively motivated language modeling research have looked into the learnability and learning efficiency of language [45, 13, 20]. By incorporating non-linguistic inputs such as multimodal stimuli [58, 32, 46] and/or communicative feedback [37, 79, 31], recent studies have explored potential mechanisms that contribute to efficient language learning in (vision-)language models. Through controlled ablation studies [71], these models can serve as proof of concept to verify cognitive mechanisms that are practically effective for machines, and generate hypotheses that are theoretically possible for human learners [43, 44]. In a similar spirit, we seek to investigate the role of explicit corrective feedback in neural language learning through controlled computational experiments. Our goal is to determine if student trials and teacher demonstrations are effective in efficient word learning in training language models.
We introduce Trial-and-Demonstration (TnD), an interactive learning framework that incorporates corrective feedback with three components: student model trials, teacher model demonstrations, and a reward conditioned on the training trajectory of the model (Figure 1). Our experiments reveal that the TnD approach accelerates language acquisition, highlighting the significance of both trials and demonstrations. From the teacher’s perspective, their word choices affect students’ word-specific learning efficiency. From the student’s side, the word frequency in trials closely aligns with the learning curve, supporting the idea that learning by language production accelerates word proficiency. Our findings highlight that trials and demonstrations can facilitate word learning in language models, and further, suggest an efficient alternative to building language models interactively.
@@ 2 Interactive Language Learning by Trials-and-Demonstrations (TnD)
Modeling corrective feedback in computational language learning presents significant challenges, as recruiting human subjects to supervise the development of a language model from the ground up over numerous iterations is impractical. Consequently, we present a human-free Trial-and-Demonstration (TnD) learning framework that streamlines the process (Figure 1). In a scenario of corrective feedback, the student model engages in production-based learning: to produce an initial utterance, followed by the teacher model generating its version of the text as a demonstration. For the student model to recognize the teacher’s response as preferable and to facilitate learning, these language outputs are evaluated by a reward function, which is based on the competence of the student’s language use that is expected for its developmental stage (i.e., training steps). The TnD framework thus includes three components: a student model’s trials, a teacher model’s demonstrations, and rewards. This framework allows us to incorporate massive corpora to study modern transformer-based generative language models, offering a general and unrestricted approach to simulate interactive language learning with corrective feedback on massive corpora.
@@ 2.1 The student model and trials
We employ randomly initialized GPT-2 [48] as the student model for our investigation into language acquisition, leveraging its causal language modeling (CLM) objective and inherent generative capabilities for production-based learning. To encourage the student model to attempt text production, it is essential to provide an appropriate context. In each trial, we prompt the student with the first 5 tokens from a natural sentence, asking it to generate the continuation as a trial.
@@ 2.2 The teacher model and demonstrations
Inspired by recent work [1, 30, 50], We utilize pre-trained language models as proxies for human language teachers. Employing language models as “caregivers” for language models offers two advantages. Firstly, it eliminates the need for recruiting human participants to engage with a language model across thousands of iterations. Secondly, we can consistently control the behavior of the teacher model across experiments, and adjust the teacher’s language behaviors by modifying its decoding strategies for language generation. The process of developing a teacher model is identical to the typical language model pre-training, as shown in Figure 1(a). We adopted the same GPT-2 architecture and pre-trained the model with the CLM objective for 100k steps, with all hyper-parameters following the default setup. To generate a natural language demonstration for the student’s trial, we prompt the pre-trained teacher model with the same 5 tokens used for the student model, thereby obtaining the teacher’s completion of the sentence.
@@ 2.3 The reward and reward model
Defining an effective reward in our context is challenging due to the absence of communication games and the lack of access to large-scale human preference annotations. Heuristic reward metrics do not consider the developmental trajectory of language models, which is critical for simulating language acquisition. It’s akin to human development where early words are celebrated as milestones, but prolonged reliance on initial language abilities can become a concern. We treat the number of training steps as the neural model’s “age” [13]. A language model that generates fluent text at 500 steps, which typically emerges around 5,000 steps, should be rewarded for its accelerated learning. Conversely, if the language production quality in the student remains the same at 50,000 steps, it should be penalized, as shown in Figure 1(c).
To train a neural age predictor, we utilize the developmental trajectory of the teacher model by saving over 100 checkpoints at various training steps. For each checkpoint, we select 25,000 contexts (each consisting of 5 tokens) from the test set and prompt the model to generate continuations for each context. This process generates a dataset comprising over 2.5 million (text, step) pairs. We then use this dataset to fine-tune a LLaMA-2-7B language model [65], incorporating a 1-layer linear head for regression. At step $n$, the student model $\pi_{\theta}$ parameterized by $\theta$ produces a sentence $\mathcal{S}$ of $l$ tokens $t_{1},\cdots,t_{l}$. We use the neural age predictor $R_{\phi}$ to estimate the logarithm of the expected training step $\hat{n}$ where this sentence typically emerges. Hence, the age-conditioned reward $r(\mathcal{S},n)$ is given by:
@@
14:         Policy update $\nabla_{\theta}\ \mathcal{L}^{\text{ppo}}_{\theta}(B)$ (See A.2)
@@ 2.4 Alternating interactive and non-interactive language learning
Reinforcement learning has been applied to language models, such as the success of games [36, 25], heuristic scores [49, 37], and models of human preferences [81, 40]. The key is to view the language production in language models as actions within a vocabulary-defined action space. Formally, the model $\pi_{\theta}$ is given a context of 5 tokens (i.e., the initial state $s_{0}=\{t_{1},\cdots,t_{5}\}$) to produce the next token (i.e., the next action $a_{1}=t_{6}$). It lands in the next state $s_{1}=\{t_{1}\cdots t_{6}\}$, and this process repeats until the sentence concludes. Following this, rewards can be computed according to Eq. 1 for the student’s trials and the teacher’s demonstrations. The goal is to maximize the expected return (i.e., the expected cumulative future reward) following $\pi_{\theta}$ along the interaction, with the trials and demonstrations both in the training batch. Inspired and taking the computational infrastructure in recent advances in reinforcement learning from human feedback (RLHF) [40], we use Proximal Policy Optimization (PPO) [54] algorithm with a clipped surrogate objective $\mathcal{L}^{\textrm{ppo}}_{\theta}$, which is the primary variant in modern large language models. [^1] We use the TRL library to train language models with reinforcement learning. We applied two modifications in the implementation of the RLHF algorithm, which include the involvement of demonstration in policy update, and the removal of KL-divergence. We refer to Appendix A.2 for mathematical details.
@@
The original PPO algorithm only learns from the agent’s own trials, i.e., the sentences it generated before. We expand the training batch to add teacher’s demonstrations: At each step, we sample the sentences generated by the latest student model and also collect those from the teacher model with the same prompts. Subsequently, we regard both of them as the training batch, and apply the policy update to improve the student model. Intuitively, our goal is to encourage the student model to imitate and repeat the teacher’s demonstration during training. Through the reward disparity between their own trial and the demonstration, we motivate the student model to learn a better language generation policy.
@@
While our interactive language learning replicates active engagement with language through corrective feedback, it’s also crucial to simulate human’s passive exposure to language, as emphasized in prior psychological [59] and computational [37] studies. We implement this with the causal language modeling objective, which is adopted by most generative language models. Consider a sentence $t_{1},\cdots,t_{l}$ in the corpus with $l$ tokens. The causal language modeling objective is to predict the next token $t_{i+1}$ given the previous context $t_{\leq i}$ by minimizing:
@@
Following the setup of Nikolaus and Fourtassi [37], we adopt an alternating learning schedule over interactive and non-interactive language learning, i.e., perform $c$ steps of causal language modeling, followed by $r$ steps of reinforcement learning, in a continuous cycle (Algorithm 1). We select $c=3$ and $r=1$, respectively, leading to 1 PPO update following every 3 CLM updates. We justify this design choice and present experiments on other hyper-parameters in Appendix B.5.
@@ 3.1 Experiment setup
We repeat our study on two training corpora: (1) the BookCorpus [80], which is commonly used for training neural language models; and (2) the BabyLM Corpus, a more developmentally plausible dataset provided by the BabyLM Challenge (Round 1) [72]. Notably, the BabyLM dataset has less than 100M words and contains a higher proportion of transcribed speech, e.g., the CHILDES corpus [33]. For each corpus, we keep 80% for training and 20% for evaluation.
@@
To reliably assess the importance or impact of a mechanism in a language learning system, computational experiments should be conducted under controlled ablation studies [71]. We describe our baselines below and study other possible setups in Appendix B.4.
@@ 3.2 Evaluation
  • The CDI set, consisting of words from the MacArthur-Bates Communicative Development Inventories (CDIs) [22]. Following Portelance et al. [45], we excluded any items comprising multiple words (such as “choo-choo”) from our dataset, as the tokenizer would treat these as distinct items.
We select these two vocabulary sets as CMN offers broader coverage while CDI is more frequency-neutral and is used to assess children’s early vocabulary development. Following Chang and Bergen [13], we remove words with less than 100 occurrences in the evaluation set of each corpus to ensure reliable results, and keep at most 512 samples for each word. This leads to 309 common words in CMN, as well as 345 words in CDI for BookCorpus and 243 words in CDI for BabyLM Corpus.
(a) Plateau v.s. unigram surprisal in BookCorpus.
(b) Plateau v.s. unigram surprisal in BabyLM.
Figure 3: Sigmoid plateau v.s. unigram surprisal.
@@
In line with previous work, we use the mean surprisal (log-perplexity) of a word to quantify the quality of the model’s predictions for this word. For each occurrence of a word $w$, the surprisal is given by $-\log_{2}P(w)$, and we average all occurrences. To visualize word acquisition from high surprisal to lower surprisal, we evaluate the student model throughout training and plot the surprisal values over logarithm training steps, leading to a learning curve for individual words and the overall vocabulary. We observe a similar pattern reported by Chang and Bergen [13] that learning curves tend to level off at a local plateau, which aligns with the unigram surprisal. This phenomenon renders the single-sigmoid model unreliable for capturing the complexity, and we adopt a double-sigmoid function to fit the learning curve. We run regression between the plateau of double-sigmoid curves and the unigram surprisals calculated from all sources of word occurrences (Figure 3). We find a strong correspondence between the plateau and unigram frequency, suggesting that the double-sigmoid function is a better option than the single-sigmoid function to fit learning curves. To analyze longer learning curves of up to 1M steps, more complex functions such as linear GAMs have been adopted [14]. For our purposes, a double-sigmoid function suffices.
Neural age of acquisition (nAoA).
To evaluate the speed at which the student model acquires a word, we employ the neural age of acquisition (nAoA). Prior research [13, 14] has used a surprisal cutoff of 50% between the minimum and maximum surprisal levels. This is akin to the method used to determine children’s age of acquisition, where the cutoff is set at the point when 50% of children are observed to produce a word [10]. To further enhance the robustness of this metric, we average nAoA over different surprisal thresholds from 0.5 to 0.95 with a step of 0.05, denoted as nAoA@[0.50:0.95]. It’s important to note that nAoA serves as a metric to assess the speed at which a model acquires a word, rather than the quality of word learning. A model might learn a word quickly, indicated by a low nAoA, yet not master it effectively, indicated by a high surprisal. Combining both metrics offers a thorough evaluation of word learning.
(a) Learning curve of CMN words on BookCorpus.
(b) Learning curve of CDI words on BookCorpus.
(c) Learning curve of CMN words on BabyLM.
(d) Learning curve of CDI words on BabyLM.
(a) Effective vocabulary size in CMN words on BookCorpus.
(b) Effective vocabulary size in CDI words on BookCorpus.
(c) Effective vocabulary size in CMN words on BabyLM.
(d) Effective vocabulary size in CDI words on BabyLM.
(a) nAoA of CMN words on BookCorpus.
(b) nAoA of CDI words on BookCorpus.
(c) nAoA of CMN words on BabyLM.
(d) nAoA of CDI words on BabyLM.
Effective vocabulary size.
[…]
@@
4 Main Results and Findings
4.1 Corrective feedback accelerates neural word acquisition
To evaluate CMN and CDI words on two corpora, we aggregate 5 random seeds and present the learning curves in Figure 4, the neural age of acquisition in Figure 6, and the effective vocabulary size in Figure 5. Figure 4 reveals that the TnD learning framework significantly accelerates word acquisition in training, outperforming other baselines. This acceleration is attributed to the critical roles of both trials and demonstrations in the learning process. With only teacher demonstrations, the student model acquires words faster than with the plain CLM baseline alone, though not as rapidly as when active trials are incorporated in the TnD framework. Conversely, without the teacher’s demonstrations, the student’s trials in the wild do not yield a marked improvement, resulting in performance comparable to the CLM baseline. We refer to Figure 14 and 15 in the Appendix for the ridgeline and scatter plot of words and their nAoA.
Figure 6 and Table 1 present nAoA at different surprisal thresholds from 0.50 to 0.95 with a step of 0.05. We find that the TnD learning framework is particularly beneficial during the earlier stages of word acquisition, as it significantly outperforms the CLM baseline on nAoA@0.50, but is eventually on-par with the CLM baseline on nAoA@0.95 towards the end of training. As a result, it can be seen from Figure 5 that students under the TnD framework quickly picked up a large volume of effective vocabulary, but eventually their vocabulary capacities have converged to the CLM baseline as expected. We further evaluate the final model on downstream natural language understanding (NLU) tasks. The TnD model performs on par with the CLM model (See Appendix B.3). Overall, our findings show that corrective feedback accelerates the student model’s neural word acquisition process, yet the student eventually converges to the teacher model’s performance. This contradicts Nikolaus and Fourtassi [37]’s previous observation that combining production and perception-based language learning from the start will deteriorate performance. We hypothesize that this discrepancy may result from using the BLEU score as a proxy for rewards from communicative feedback in their work, while our setup incorporates a more explicit form of corrective feedback.
@@ 4.1 Corrective feedback accelerates neural word acquisition
Figure 7: On 2 training corpora and 2 test vocabulary, we further present the fitted learning curves on smaller student models ($\texttt{TnD}_{d=588/360/250}$) of mean surprisal over $\log_{10}$ training steps, with nAoA@0.5 of each curve indicated by a vertical dashed line.
@@ 4.2 Corrective feedback helps knowledge distillation for smaller student models
The current age-based reward design assumes the student and teacher models are of the same size. This section investigates if such a reward function could distill linguistic knowledge to smaller student models. The original student GPT-2 model has a dimension of $d=768$ (12 attention heads each with a dimension of 64). We now keep all experimental setups untouched but smaller student models with dimensions of 588 ($12\times 49$), 360 ($10\times 36$), and 250 ($10\times 25$) respectively. Figure 7 shows that the efficiency in early language learning can still be observed. Each TnD model outperforms the CLM baseline of the same size, and even surpasses CLM baselines of large capacity in early steps.
(a) The word “have.”
(b) The word “now.”
(c) The word “see.”
(d) The word “again.”

### v2

@@ Abstract
Humans are efficient language learners and inherently social creatures. Our language development is largely shaped by our social interactions, for example, the demonstration and feedback from caregivers. Contrary to human language learning, recent advancements in large language models have primarily adopted a non-interactive training paradigm, and refined pre-trained models through feedback afterward. In this work, we explore how corrective feedback from interactions influences neural language acquisition from scratch through systematically controlled experiments, assessing whether it contributes to word learning efficiency in language models. We introduce a trial-and-demonstration (TnD) learning framework that incorporates three distinct components: student trials, teacher demonstrations, and a reward conditioned on language competence at various developmental stages. Our experiments reveal that the TnD approach accelerates word acquisition for student models of equal and smaller numbers of parameters, and we highlight the significance of both trials and demonstrations. We further show that the teacher’s choices of words influence students’ word-specific learning efficiency, and a practice-makes-perfect effect is evident by a strong correlation between the frequency of words in trials and their respective learning curves. Our findings suggest that interactive language learning, with teacher demonstrations and active trials, can facilitate efficient word learning in language models.
Babysit A Language Model From Scratch:
Interactive Language Learning by Trials and Demonstrations
Ziqiao Mathanks: Authors contributed equally to this work. University of Michigan marstin@umich.edu                        Zekun Wang11footnotemark: 1 University of Michigan zekun@umich.edu                        Joyce Chai University of Michigan chaijy@umich.edu
@@ 1 Introduction
Humans are social beings and we learn language from interactions (Vygotsky, 1934; Bruner, 1985; Palincsar, 1986; Kuhl, 2004; Tomasello, 2005). Long before children’s linguistic skills are mature, they could engage in early forms of conversational exchange with others (Halliday, 1975; Clark, 2018). A critical component of social interactions that language grounds to is the feedback provided by the caregivers (Warlaumont et al., 2014). This, for example, includes the communicative feedback (Bates et al., 1975; Nikolaus and Fourtassi, 2023) that highlights the success and failure of communication, and the corrective feedback (Farrar, 1992; Chouinard and Clark, 2003; Saxton et al., 2005; Hiller, 2016) that is more direct and emphasizes the responses from caregivers, which offer corrections to possible errors in children’s speech, encompassing variants like negative evidence, reformulations, or recasts. Unlike human learners who acquire language skills through feedback during interactions, most language models differ in terms of their inductive biases and data sources (Warstadt and Bowman, 2022). These models typically learn from massive text corpora using cross-entropy loss for self-supervised learning.
@@
Recently, several lines of cognitively motivated language modeling research have looked into the learnability and learning efficiency of language Portelance et al. (2020); Chang and Bergen (2022); Evanson et al. (2023). By incorporating non-linguistic inputs such as multimodal stimuli Shi et al. (2019); Ma et al. (2023); Portelance et al. (2024) and/or communicative feedback Nikolaus and Fourtassi (2021); Zhu et al. (2022); Liu et al. (2022), recent studies have explored potential mechanisms that contribute to efficient language learning in (vision-)language models. Through controlled ablation studies Warstadt and Bowman (2022), these models can serve as proof of concept to verify mechanisms that are practically effective for machines, and generate hypotheses that are possible for cognitive learners (Portelance, 2022; Portelance and Jasbi, 2023). In a similar spirit, we seek to investigate the role of explicit corrective feedback in neural language learning through controlled computational experiments. Rather than making direct comparisons to human learning, our goal is to examine if student trials and teacher demonstrations promote efficient word learning in language model training, and if so, which words benefit the most.
We introduce Trial-and-Demonstration (TnD), an interactive learning framework that incorporates corrective feedback with three components: student model trials, teacher model demonstrations, and a reward conditioned on the training trajectory of the model (Figure 1). Our experiments reveal that the TnD approach accelerates word acquisition, highlighting the significance of both trials and demonstrations. From the teacher’s perspective, their word choices affect students’ word-specific learning efficiency. From the student’s side, the word frequency in trials closely aligns with the learning curve, supporting the idea that learning by language production accelerates word proficiency. Our findings highlight that trials and demonstrations can facilitate word learning in language models, and further, suggest an efficient alternative to building language models interactively.
@@ 2 Interactive Language Learning by Trials-and-Demonstrations (TnD)
Modeling corrective feedback in computational language learning presents significant challenges, as recruiting human subjects to supervise the development of a language model from the ground up over numerous iterations is impractical. Consequently, we present a human-free Trial-and-Demonstration (TnD) learning framework that streamlines the process (Figure 1). In a scenario of corrective feedback, the student model engages in production-based learning: to produce an initial utterance, followed by the teacher model generating its version of the text as a demonstration. For the student model to recognize the teacher’s response as preferable and to facilitate learning, these language outputs are evaluated by a reward function, which is based on the competence of the student’s language use that is expected for its developmental stage (i.e., training steps). The TnD framework thus includes three components: a student model’s trials, a teacher model’s demonstrations, and rewards. This framework allows us to incorporate massive corpora to study modern generative language models, offering a general and unrestricted approach to simulate interactive language learning with corrective feedback on massive corpora.
@@ 2.1 The student model and trials
We employ randomly initialized GPT-2 (Radford et al., 2019) as the student model for our investigation into language acquisition, leveraging its causal language modeling (CLM) objective and inherent generative capabilities for production-based learning. To encourage the student model to attempt text production, it is essential to provide an appropriate context. In each trial, we prompt the student with the first 5 tokens from a natural sentence, asking it to generate the continuation as a trial.
@@ 2.2 The teacher model and demonstrations
Inspired by recent work (Bai et al., 2022; Lee et al., 2023; Saha et al., 2023), we utilize pre-trained language models as proxies for human language teachers. Employing language models as “caregivers” for language models offers two advantages. Firstly, it eliminates the need for recruiting human participants to engage with a language model across thousands of iterations. Secondly, we can consistently control the behavior of the teacher model across experiments, and adjust the teacher’s language behaviors by modifying its decoding strategies for language generation. The process of developing a teacher model is identical to the typical language model pre-training, as shown in Figure 1(a). We adopted the same GPT-2 architecture and pre-trained the model with the CLM objective for 100k steps, with all hyper-parameters following the default setup. To generate a natural language demonstration for the student’s trial, we prompt the pre-trained teacher model with the same 5 tokens used for the student model, thereby obtaining the teacher’s completion of the sentence.
@@ 2.3 The reward and reward model
Defining an effective reward in our context is challenging due to the absence of communication games and the lack of access to large-scale human preference annotations. Heuristic reward metrics do not consider the developmental trajectory of language models, which is critical for simulating language acquisition. It’s akin to human development where early words are celebrated as milestones, but prolonged reliance on initial language abilities can become a concern. We treat the number of training steps as the neural model’s “age” (Chang and Bergen, 2022). A language model that generates fluent text at 500 steps, which typically emerges around 5,000 steps, should be rewarded for its accelerated learning. Conversely, if the language production quality in the student remains the same at 50,000 steps, it should be penalized. This process is illustrated in Figure 1(c).
To train a neural age predictor, we utilize the developmental trajectory of the teacher model by saving over 100 checkpoints at various training steps. For each checkpoint, we select 25,000 contexts (each consisting of 5 tokens) from the test set and prompt the model to generate continuations for each context. This process generates a dataset comprising over 2.5 million (text, step) pairs. We then use this dataset to fine-tune a LLaMA-2-7B language model (Touvron et al., 2023), incorporating a 1-layer linear head for regression. At step $n$, the student model $\pi_{\theta}$ parameterized by $\theta$ produces a sentence $\mathcal{S}$ of $l$ tokens $t_{1},\cdots,t_{l}$. We use the neural age predictor $R_{\phi}$ to estimate the logarithm of the expected training step $\hat{n}$ where this sentence typically emerges. Hence, the age-conditioned reward $r(\mathcal{S},n)$ is given by:
@@
14:         Policy update $\nabla_{\theta}\ \mathcal{L}^{\text{ppo}}_{\theta}(B)$ (See A.3)
@@ 2.4 Alternating interactive and non-interactive language learning
Reinforcement learning has been applied to language models, such as the success of games (Narasimhan et al., 2015; He et al., 2016), heuristic scores (Ranzato et al., 2016; Nikolaus and Fourtassi, 2021), and models of human preferences (Ziegler et al., 2019; Ouyang et al., 2022). The key is to view the language production in language models as actions within a vocabulary-defined action space. Formally, the model $\pi_{\theta}$ is given a context of 5 tokens (i.e., the initial state $s_{0}=\{t_{1},\cdots,t_{5}\}$) to produce the next token (i.e., the next action $a_{1}=t_{6}$). It lands in the next state $s_{1}=\{t_{1}\cdots t_{6}\}$, and this process repeats until the sentence concludes. Following this, rewards can be computed according to Eq. 1 for the student’s trials and the teacher’s demonstrations. The goal is to maximize the expected return (i.e., the expected cumulative future reward) following $\pi_{\theta}$ along the interaction, with the trials and demonstrations both in the training batch. Inspired and taking the computational infrastructure in recent advances in reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), we use Proximal Policy Optimization (PPO) Schulman et al. (2017) algorithm with a clipped surrogate objective $\mathcal{L}^{\textrm{ppo}}_{\theta}$, which is the primary variant in modern large language models. [^1] We use the TRL library to train language models with reinforcement learning. Our code is available here. We applied two modifications in the implementation of the RLHF algorithm, which include the involvement of demonstration in policy update, and the removal of KL-divergence. We refer to Appendix A.3 for mathematical details.
@@
The original PPO algorithm only learns from the agent’s own trials, i.e., the sentences it generated before. We expand the training batch to add teacher’s demonstrations: At each step, we sample the sentences generated by the latest student model and also collect those from the teacher model with the same prompts. Subsequently, we regard both of them as the training batch, and apply the policy update to improve the student model. Intuitively, our goal is to encourage the student model to imitate and repeat the teacher’s demonstration during training. As there is a reward disparity between the student’s trial and the teacher’s demonstration, we motivate the student model to learn a better language generation policy.
@@
While our interactive language learning replicates active engagement with language through corrective feedback, it’s also crucial to simulate human’s passive exposure to language, as emphasized in prior psychological (Smith and Yu, 2008) and computational (Nikolaus and Fourtassi, 2021) studies. We implement this with the causal language modeling objective, which is adopted by most generative language models. Consider a sentence $t_{1},\cdots,t_{l}$ in the corpus with $l$ tokens. The causal language modeling objective is to predict the next token $t_{i+1}$ given the previous context $t_{\leq i}$ by minimizing:
@@
Following the setup of Nikolaus and Fourtassi (2021), we adopt an alternating learning schedule over interactive and non-interactive language learning, i.e., perform $c$ steps of causal language modeling, followed by $r$ steps of reinforcement learning, in a continuous cycle (Algorithm 1). We select $c=3$ and $r=1$, respectively, leading to 1 PPO update following every 3 CLM updates. We justify this design choice and present experiments on other hyper-parameters in Appendix B.6.
@@ 3 Experiment and Evaluations
(a) Learning curve of CMN words on BookCorpus.
(b) Learning curve of CDI words on BookCorpus.
(c) Learning curve of CMN words on BabyLM.
(d) Learning curve of CDI words on BabyLM.
(a) nAoA of CMN words on BookCorpus.
(b) nAoA of CDI words on BookCorpus.
(c) nAoA of CMN words on BabyLM.
(d) nAoA of CDI words on BabyLM.
(a) Effective vocabulary size in CMN words on BookCorpus.
(b) Effective vocabulary size in CDI words on BookCorpus.
(c) Effective vocabulary size in CMN words on BabyLM.
(d) Effective vocabulary size in CDI words on BabyLM.
@@ 3.1 Experiment setup
We repeat our study on two training corpora: (1) the BookCorpus (Zhu et al., 2015), which is commonly used for training neural language models; and (2) the BabyLM Corpus, a more developmentally plausible dataset provided by the BabyLM Challenge (Round 1) (Warstadt et al., 2023). Notably, the BabyLM dataset has less than 100M words and contains a higher proportion of transcribed speech, e.g., the CHILDES corpus (MacWhinney, 2000). For each corpus, we keep 80% for training and 20% for evaluation.
@@
To reliably assess the importance or impact of a mechanism in a language learning system, computational experiments should be conducted under controlled ablation studies (Warstadt and Bowman, 2022). We describe our baselines below and study other possible setups in Appendix B.5.
@@ 3.2 Evaluation
  • The CDI set, consisting of words from the MacArthur-Bates Communicative Development Inventories (CDIs) (Fenson et al., 2006). Following Portelance et al. (2020), we excluded any items comprising multiple words (such as “choo-choo”) from our dataset, as the tokenizer would treat these as distinct items.
We select these two vocabulary sets as CMN offers broader coverage while CDI is more frequency-neutral and is used to assess children’s early vocabulary development. Following Chang and Bergen (2022), we remove words with less than 100 occurrences in the evaluation set of each corpus to ensure reliable results, and keep at most 512 samples for each word. This leads to 309 common words in CMN, as well as 345 words in CDI for BookCorpus and 243 words in CDI for BabyLM Corpus.
@@
In line with previous work, we use the mean surprisal (log-perplexity) of a word to quantify the quality of the model’s predictions for this word. For each occurrence of a word $w$, the surprisal is given by $-\log_{2}P(w)$, and we average all occurrences. To visualize word acquisition from high surprisal to lower surprisal, we evaluate the student model throughout training and plot the surprisal values over logarithm training steps, leading to a learning curve for individual words and the overall vocabulary. We observe a similar pattern reported by Chang and Bergen (2022) that learning curves tend to level off at a local plateau, which aligns with the unigram surprisal. This phenomenon renders the single-sigmoid model unreliable for capturing the complexity, and we adopt a double-sigmoid function to fit the learning curve. We justify this choice and discuss more on the patterns in Appendix B.2.
@@
Neural age of acquisition (nAoA).
To evaluate the speed at which the student model acquires a word, we employ the neural age of acquisition (nAoA). Prior research (Chang and Bergen, 2022; Chang et al., 2023) has used a surprisal cutoff of 50% between the minimum and maximum surprisal levels. This is akin to the method used to determine children’s age of acquisition, where the cutoff is set at the point when 50% of children are observed to produce a word (Braginsky et al., 2016). To further enhance the robustness of this metric, we average nAoA over different surprisal thresholds from 0.5 to 0.95 with a step of 0.05, denoted as nAoA@[0.50:0.95]. It’s important to note that nAoA serves as a metric to assess the speed at which a model acquires a word, rather than the quality of word learning. A model might learn a word quickly, indicated by a low nAoA, yet not master it effectively, indicated by a high surprisal. Combining both metrics offers a thorough evaluation of word learning.
@@
Figure 6: We further present the fitted learning curves on smaller student models ($\texttt{TnD}_{d=588/360/250}$) of mean surprisal over $\log_{10}$ training steps, with nAoA@0.5 of each curve indicated by a vertical dashed line.
@@
Effective vocabulary size.
[…]
4 Main Results and Findings
4.1 Corrective feedback accelerates neural word acquisition
To evaluate CMN and CDI words on two corpora, we aggregate 5 random seeds and present the learning curves in Figure 3, the neural age of acquisition in Figure 4, and the effective vocabulary size in Figure 5. Figure 3 reveals that the TnD learning framework significantly accelerates word acquisition in training, outperforming other baselines. This acceleration is attributed to the critical roles of both trials and demonstrations in the learning process. With only teacher demonstrations, the student model acquires words faster than with the plain CLM baseline alone, though not as rapidly as when active trials are incorporated in the TnD framework. Conversely, without the teacher’s demonstrations, the student’s trials in the wild do not yield a marked improvement, resulting in performance comparable to the CLM baseline. We refer to Figure 13 and 14 in the Appendix for the ridgeline and scatter plot of words and their nAoA.
Figure 4 and Table 1 present nAoA at different surprisal thresholds from 0.50 to 0.95 with a step of 0.05. We find that the TnD learning framework is particularly beneficial during the earlier stages of word acquisition, as it significantly outperforms the CLM baseline on nAoA@0.50, but is eventually on-par with the CLM baseline on nAoA@0.95 towards the end of training. As a result, it can be seen from Figure 5 that students under the TnD framework quickly picked up a large volume of effective vocabulary, but eventually their vocabulary capacities have converged to the CLM baseline as expected. We further evaluate the final model on downstream natural language understanding (NLU) tasks. The TnD model performs on par with the CLM model (See Appendix B.4). Overall, our findings show that corrective feedback accelerates the student model’s neural word acquisition process, yet the student eventually converges to the teacher model’s performance. This contradicts the previous observations by Nikolaus and Fourtassi (2021) that combining production and perception-based language learning from the start will deteriorate performance. We hypothesize that this discrepancy may result from using the BLEU score as a proxy for rewards from communicative feedback in their work, while our setup incorporates a more explicit form of corrective feedback.